OpenStudy (anonymous):

What graph goes with these following equations? A)2x – 2y = 8 4x + 4y = 16 B)2x – 2y = 8 4x – 4y = 16 C)2x + 2y = 8 4x + 4y = 16 D)2x + 2y = 8 4x – 4y = 16

OpenStudy (anonymous):

here is the graph

hero (hero):

There's no intersection

hero (hero):

Hmmm

hero (hero):

That's because it's the same graph. Now we just have to figure out which one is right. in this case we have to use the x and y intercepts to figure this out

OpenStudy (anonymous):

c..

OpenStudy (anonymous):

x/4=y/4=1 this show that in both axis the cutten part is 4,0 and 0,4

OpenStudy (anonymous):

so its B? are you sure?

hero (hero):

Yeah, I realize that for c, the slope is negative so it's c

OpenStudy (anonymous):

C) The equation for the line on the plot is x + y = 4 All of the other alternatives offered have sign problems.